Use factor trees to find the GCF and LCM of the following:

a. 24 and 60 b. 150, 210 and 375


1
Expert's answer
2022-03-17T15:46:13-0400

A) 24 and 60

24 and 60

4×6 and 4×15

2×2×2×3 and 2×2×3×5

GCF will be ( 2×2×3= 12)

LCM will be ( 2×2×3×5= 60)


B) 150

2×3×5×5


210

2×3×5×7


375

3×5×5×5

GCF will be ( 3×5= 15)

LCM will be 2×3×5×5×5×7= 5250